OpenStudy (anonymous):

With the equation of a line through the given points or through the given point with the given slope (Algebra I)? (5,7), (6,8)

OpenStudy (anonymous):

**write

OpenStudy (anonymous):

Both slope interject and point slope form.

OpenStudy (anonymous):

intercept**

OpenStudy (anonymous):

you got the slope?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

you know how to find it?

OpenStudy (anonymous):

Not really :/ rate of change??

OpenStudy (anonymous):

yeah first we can do it in our heads, then we can do it the donkey way from 5 to 6 is one unit right in the x direction from 7 to 8 is one unit up in the y direction right one, up one slope is one

OpenStudy (anonymous):

so the slope is 1

OpenStudy (anonymous):

else we compute \[\frac{y_2-y_1}{x_2-x_1}=\frac{8-7}{6-5}=\frac{1}{1}=1\]

OpenStudy (anonymous):

yeah it is \(1\) now use \[y-y_1=m(x-x_1)\] with either point as \((x_1,y_1)\) and \(m=1\)

OpenStudy (anonymous):

kk can you finish it with me?

OpenStudy (anonymous):

plug one in to the equation y-y1=m(x-x1)

OpenStudy (anonymous):

sure we will put the first pair of numbers in the point slope formula \[y-7=1(x-5)\] of course we don't need the 1, so it is really \[y-7=x-5\]

OpenStudy (anonymous):

solve for \(y\) by adding 7 to both sides

OpenStudy (anonymous):

for y1 and x1 use any one of the cordinates